# Fisher-Orthogonal Projected Natural Gradient Descent for Continual Learning

### v1

@@ Appendix B Hyperparameter Values
For the learning rate, we swept over [1e-5, 5e-5, 1e-4, 5e-4, 1e-3, 5e-3, 1e-2, 5e-2, 1e-1]. For EWC, the $\lambda$ regularization coefficient was swept over [10, 50, 100, 400]. For FOPNG, FNG, FOPNG-PreFisher, the $\lambda$ regularization coefficient was swept over [1e-4, 5e-4, 1e-3, 1e-2] after choosing the optimal learning rate while using $\lambda=$1e-3; it is possible that a more thorough sweep over FOPNG’s hyper-parameters could yield slightly higher performance. Section C shows that performance is extremely stable for different FOPNG $\alpha$ values, so we fix $\alpha=0.5$ for all experiments.

### v2

@@ Appendix B Hyperparameter Values
For the learning rate, we swept over [1e-5, 5e-5, 1e-4, 5e-4, 1e-3, 5e-3, 1e-2, 5e-2, 1e-1]. For EWC, the $\lambda$ regularization coefficient was swept over [10, 50, 100, 400]. For FOPNG, FNG, FOPNG-PreFisher, the $\lambda$ regularization coefficient was swept over [1e-4, 5e-4, 1e-3, 1e-2] after choosing the optimal learning rate while using $\lambda=$1e-3; it is likely that a more thorough sweep over FOPNG’s hyper-parameters would yield slightly higher performance. Section C shows that performance is extremely stable for different FOPNG $\alpha$ values, so we fix $\alpha=0.5$ for all experiments.